OpenStudy (anonymous):

For the simple harmonic motion equation d=5sin((pi/4)t), what is the period?

OpenStudy (anonymous):

Let WolframAudio help you: http://www.wolframalpha.com Enter: d=5sin(pi/4t) Solution: http://www.wolframalpha.com/input/?i=d%3 … Look at the plots. The variable t (period) is: t = pi/(4sin^(−1)(d/5))

OpenStudy (anonymous):

w = pi /4 = 2 pi f 1 /4 = 2 f f = 1/8 = angular frequency = 1 / period period = 8 unit

OpenStudy (anonymous):

period is :after what time will the body return to same phase..after what consecutive time interval will the body be at d=5...it will be at d=5 at t=2 (check it) and t=10(check it too!!)...so diff between them is 8 sec and so is the period.. else use the aforesaid formulas..